# MultiCrafter: High-Fidelity Multi-Subject Generation via Disentangled Attention and Identity-Aware Preference Alignment

### v1

@@ 4.4 Identity-Preserving Preference Optimization
To further enhance the generation quality and align with human preferences, we introduce a post-training stage using reinforcement learning. This final stage aims to refine aesthetic appeal and text-image alignment without compromising the subject fidelity. We adapt the efficient MixGRPO framework (Li et al., 2025b), which confines stochastic optimization to a sliding window $S$. However, standard GRPO with its token-level policy ratios can exhibit instability, particularly when training MoE models due to expert routing fluctuations (Zheng et al., 2025). To mitigate this and better suit our MoE-LoRA architecture from Sec.˜4.3, we replace the GRPO objective with the more stable Group Sequence Policy Optimization (GSPO) (Zheng et al., 2025). Specifically, we replace the token-level policy ratio in Eq.˜3 with a sequence-level policy ratio $s_{i}(\theta)$ defined over the denoising steps within the sliding window $S$:
@@
This sequence-level ratio reflects the overall policy shift for the entire sequence within the optimization window, leading to more stable gradients. The advantage score $A_{i}$ is calculated as in Eq.˜3, but using our composite reward from Eq.˜9. The final optimization objective is thus:
$J_{\text{MixGSPO}}(\theta)=\mathbb{E}_{c,\mathcal{Z},\{x_{i}^{T}\}_{i=0}^{N}\sim\pi_{\theta_{\text{old}}}(\cdot|c,\mathcal{Z})}\left[\frac{1}{N}\sum_{i=1}^{N}\min\Big(s_{i}(\theta)A_{i},\text{clip}(s_{i}(\theta),1-\beta,1+\beta)A_{i}\Big)\right].$ (8)
@@
where $R_{text}$ is the text alignment reward from a pre-trained CLIP model, $R_{aes}$ is an aesthetic reward from a predictor like HPSv2 (Wu et al., 2023), $R_{id}$ is used to evaluate subject fidelity, and $w_{id},w_{text},w_{aes}$ are their corresponding weights, . To accurately measure the subject fidelity of multi-subject generation results, we built the Multi-ID Alignment Reward using the Hungarian matching algorithm. For human subjects, we first employ a face detector (Deng et al., 2020) to extract facial embeddings from each reference image. We then apply the same detector to the generated image to identify all present faces and extract their embeddings. Then we construct a pairwise similarity matrix $C$ where $C_{ij}$ is the cosine similarity between the embedding of the $i$-th reference face and the $j$-th detected face. The Hungarian algorithm is then used to solve the assignment problem by finding an assignment matrix $X\in\{0,1\}^{N_{ref}\times N_{gen}}$ that maximizes the total similarity:
@@
where $N_{ref}$ and $N_{gen}$ are the number of reference and generated faces, respectively. These ensure each face is matched at most once, preventing reward hacking, stopping the model from using attribute leakage to generate multiple “average faces" for an unearned high reward. For object subjects, each reference object is pre-annotated with a text prompt. We leverage Florence-2 (Xiao et al., 2024) and SAM2 (Ravi et al., 2024) to locate the corresponding object in the generated image. Then we compute the cosine similarity between the DINOv2 (Oquab et al., 2023) embeddings of the segmented region and the reference object. We provide the complete pseudocode at Appendix˜A.
@@ 5.1 Experimental Setup
CLIP-T Face-Sim DINO-I CLIP-I HPS CLIP-T DINO-I CLIP-I HPS AVG
@@
Table 1: Quantitative comparison with state-of-the-art methods on the multi-human and multi-object generation. The best results are in bold, and the second-best are underlined. Our method outperforms others, especially in subject fidelity metrics, and achieves the highest overall average score.
@@
Implementation Details. To achieve precise multi-subject customized generation, our training process is divided into three stages: Single-Subject Pre-training, Multi-Subject Customization Training, and Identity-Preserving Preference Optimization. Following (Wu et al., 2025c), we set the resolution of generated images to $512\times 512$ , the resolution of reference images to $320\times 320$, and each LoRA module’s rank to $r=512$. For the MoE-LoRA applied to the FFN layers, we configure it with 4 experts and activate 1 expert per forward pass. For reinforcement learning, we configure with a sampling step 16, a window size of $w=2$, a shift interval of $\tau=50$, and a window stride of $s=1$. More details settings can be found in Appendix˜B.
@@
Datasets and BenchMark. We constructed separate datasets for multi-human and multi-object generation. Due to the scarcity of public multi-human customization datasets with adequate annotations, we designed a data collection pipeline to curate a new dataset from videos for multi-human customization. We supplemented the details in the Appendix˜C. A total of 200k pairs of cross-pair data were obtained for training. For the reinforcement learning stage, we curated a face collection of 80 celebrities and 80 non-celebrities. We paired the faces in the collection and used Qwen2.5-VL (Bai et al., 2025) to generate a prompt for each pair, resulting in 12,720 data points. We randomly selected 1,000 of these data points as our benchmark, and the remaining data was used as reinforcement learning training data. For multi-object customization, we use the public MUSAR-Gen (Guo et al., 2025) Dataset as our foundation. We use Florence-2 and SAM to obtain the detailed positions of reference objects. To ensure segmentation quality, we further employ Qwen2.5-VL to filter the results. We split 1,000 samples from this dataset that were not seen during training as a benchmark.
@@
Evaluation metrics. Following the  (Le et al., 2025; Mou et al., 2025), we evaluate generated image quality using standard metrics. We calculate the cosine similarity between the prompt and the image CLIP (Radford et al., 2021) embeddings (CLIP-T) to evaluate text fidelity. We also use HPSv2 for aesthetic and human preference scoring. For subject fidelity, we employ cosine similarity measures between generated images and reference images within CLIP and DINO (Zhang et al., 2022) spaces, referred to as CLIP-I and DINO-I scores, respectively. Note that in order to accurately calculate CLIP-I and DINO-I, we first use a face detector (Deng et al., 2019) or Florence-2 to accurately locate the position of subjects in the generated image, and then calculate them. Additionally, for multi-human generation, we incorporate Face Similarity (Face-Sim) (Deng et al., 2019) and the Hungarian algorithm, enabling them to assess subject fidelity more accurately.
@@
Figure 4: Qualitative comparison with existing methods on the multi-human generation (Zoom in for best visual comparison). Our method significantly improves subject fidelity.
@@ 5.2 Quantitative Comparison
We conducted comprehensive comparisons with existing methods in both multi-subject and multi-object customized generation. These methods include MS-Diffusion (Wang et al., 2025c), MIP-Adapter (Zhong et al., 2025), OmniGen (Xiao et al., 2025), UNO (Wu et al., 2025c), OmniGen2 (Wu et al., 2025b), DreamO (Mou et al., 2025), and XVerse (Chen et al., 2025a). The results, as shown in Tab.˜1, demonstrate that our method achieves significant improvements over existing approaches, particularly in terms of subject fidelity. In the multi-human generation task, our model shows a commanding lead in identity preservation, achieving the top scores in subject fidelity. Notably, the significant 28.3% relative improvement in the Face-Sim over the next-best method highlights our model’s ability to effectively distinguish between different subjects and preserve their individual detailed features. This is crucial for accurately and reliably accomplishing the highly sensitive and challenging task of multi-human customization. Concurrently, the model remains highly competitive in text-image alignment, achieving a strong balance between subject fidelity and prompt alignment. This outstanding performance extends to the multi-object generation benchmark, where our method again secures the top ranks in text alignment and subject fidelity, validating its robustness and versatility. While HPS score tends to vibrant yet unnatural ("oily") outputs, our model is specifically trained for multi-human customization with an emphasis on photorealism. To preserve this quality, we follow Xue et al. (2025) and employ the CLIP Score during reinforcement learning. Therefore, we only achieve competitive objective scoring results in aesthetics, but we can generate more realistic results. Ultimately, by achieving the best overall score, our method achieves significant improvements over existing methods, validating the effectiveness of our framework.
@@ 5.3 Qualitative Comparison
We provide a qualitative comparison against existing methods on the more challenging task of multi-human customized generation in Fig.˜4. As shown in the first two rows of Fig.˜4, methods trained directly with In-Context Learning, such as UNO, OmniGen, and OmniGen2, struggle with attribute confusion when generating subjects of the same gender, which degrades subject fidelity. In contrast, our method accurately preserves the unique features of each individual. Our method maintains strong subject fidelity even in interactive scenarios, as shown in the third row of Fig.˜4. This result validates the effectiveness of our framework. Aesthetically, while DreamO and OmniGen2 produce vibrantly colored images, our approach generates images with a higher degree of realism. Additional qualitative comparisons for both multi-subject generation (includes both human and object) are provided in the Appendices˜G, H and I.
@@ 5.4 Ablation Studies
IDAR MoE-LoRA IPPO CLIP-T Face-Sim DINO-I CLIP-I HPS
0.2645 0.1474 0.5972 0.6489 0.2954
@@
Table 2: Ablation study of the core module of MultiCrafter. IDAR is Identity-Disentangled Attention Regularization (Sec.˜4.2), MoE-LoRA is our Efficient Adaptive Expert Tuning (Sec.˜4.3) , and IPPO is our Identity-Preserving Preference Optimization (Sec.˜4.4). The results demonstrate the effectiveness of each module.
We conduct detailed ablation studies to validate the effectiveness of each component of our proposed framework. Since the multi-human customization evaluation metrics is more accurate, we conducted an ablation experiment on multi-human customization. We use UNO as our baseline for comparison, and the detailed results are presented in Tab.˜2. The baseline model relies solely on a simple reconstruction loss. As shown in Tab.˜2, the base model performs poorly in terms of subject fidelity, achieving a Face-Sim score of only 0.1474. This confirms that a simple objective is insufficient for handling complex multi-subject scenarios. Upon introducing our Identity-Disentangled Attention Regularization (IDAR), we observe a dramatic improvement across all subject fidelity metrics. As shown on the left of Fig.˜2, our method successfully separates the attention of different subjects. This result underscores the critical role of our IDAR in explicitly disentangling subject features and preventing attribute leakage. We attribute the slight decrease in the HPS score to the quality of our training dataset, a point we discuss further in the Appendix˜E. Building on this, we integrate the MoE-LoRA architecture, which yields further gains across subject fidelity while also improving text alignment. This demonstrates that our strategy of using MoE to allow different experts to specialize in handling diverse spatial layouts is effective in enhancing generation quality. Finally, Identity-Preserving Preference Optimization (IPPO) further boosts subject fidelity. More importantly, it significantly enhances alignment with human preferences, leading to notable improvements in both text alignment and aesthetic scores. This shows that our IPPO successfully aligns the model’s output with human preferences without compromising the subject fidelity.
@@ 6 Conclusion
In this paper, we introduced MultiCrafter, a novel framework designed to address the critical challenges of feature bleed and identity degradation in multi-subject customized generation. Our framework uses Identity-Disentangled Attention Regularization to prevent "attention bleeding" and alleviate the degradation of subject fidelity caused by attribute leakage. Then, we introduce a Mixture-of-Experts architecture to enhance model capacity. We further align the model with human preferences using a novel online reinforcement learning framework featuring a Multi-ID Alignment Reward and the stable GSPO algorithm. Experiments show that MultiCrafter significantly improves subject fidelity while better aligning with human preferences and generating realistic images.
@@ Appendix A More Details for Identity-Preserving Preference Optimization
We provide a detailed algorithm execution flow for Identity-Preserving Preference Optimization Sec.˜4.4, as shown in Algorithm˜1.
@@
In this section, we provide more details on the hyperparameter settings and specific training details. As mentioned in the main text, we divide the training phase into three parts, and the details of each phase are as follows: 1. Single-Subject Pre-training. We first pre-train the model for 40,000 steps on an internal single-subject dataset to equip it with foundational subject customization capabilities. In this stage, only $\mathcal{L}_{diff}$ is used for supervision. We use the AdamW optimizer with a $3\times 10^{-5}$ learning rate and a weight decay of $1\times 10^{-2}$. We use 8 cards for training and set the batch size of each card to 6. 2. Multi-Subject Customization Training. Then we train two models for customized human generation and object generation, respectively. For multi-human customized generation, we decrease the learning rate to $1\times 10^{-5}$, set the loss weight $\lambda=0.3$, and introduce the attention loss $\mathcal{L}_{attn}$. This stage runs for 25,000 steps. We use 8 cards at this stage and set the batch size to 4. For multi-object customized generation, since the size of the dataset is smaller than multi-human datasets, we only train for 15,000 steps. 3. Identity-Preserving Preference Optimization. Finally, we fine-tune the model using our proposed reinforcement learning method. Following (Li et al., 2025b), we configure with a sampling step of 16, a window size of $w=2$, a shift interval of $\tau=50$, and a window stride of $s=1$. This stage consists of 300 steps. For multi-human customization, the reward weights are set to $w_{id}=0.5$, $w_{text}=1.4$, and $w_{aes}=0.7$. For multi-object customization, we adjust the subject fidelity weight to $w_{id}=1.0$, while keeping $w_{text}=1.4$ and $w_{aes}=0.7$. In this stage, we used 16 cards for training and set the batch size of each card to 1.
@@
Due to the scarcity of public multi-human customization datasets with adequate annotations, we designed a comprehensive and automated data curation pipeline as shown in Fig.˜5. This pipeline processes raw video clips to generate structured training samples, each containing a target image with multiple subjects, corresponding identity reference images, segmentation masks, and a detailed textual description. The entire workflow ensures subject fidelity, high image quality, and rich annotation. Specifically, we sample video clips featuring two individuals from a large database. The process begins with frame selection and subject localization. For each input video, we sample an initial frame as the source for reference images and a middle frame as the target scene. We first employ a YOLO-Pose (Maji et al., 2022) to obtain initial bounding boxes and keypoint information for each person. Following localization, we leverage the Segment Anything Model (SAM) (Ravi et al., 2024) to generate high-fidelity segmentation masks for each individual, effectively isolating them from the background. To refine this output, only the largest connected component of the mask is retained. A critical subsequent step is ensuring subject fidelity across frames. We apply a face detection modell (Deng et al., 2019) to the segmented portraits to locate facial regions, and then use a face recognition model (Deng et al., 2019) to extract a normalized feature embedding for each face. By computing the cosine similarity between embeddings from the reference and target frames, we enforce a stringent threshold to discard pairs where identity cannot be confidently verified. Once a pair of frames passes this verification, the pipeline generates the final training sample. The segmented portraits and cropped faces from the initial frame are saved as the reference =images. The target frame is cropped into a square as target image, with its corresponding body and face masks preserved. Finally, a powerful vision-language model, Qwen2.5-VL (Bai et al., 2025), is prompted with the reference images and the target image to produce a rich text prompt of the entire scene, ensuring descriptive consistency for each subject.
@@
Figure 5: Data processing pipeline for customized multi human image generation.
Appendix D More Details for Our Benchmark.

### v2

@@ 4.3 Identity-Preserving Preference Optimization
To further enhance the generation quality and align with human preferences, we introduce a post-training stage using reinforcement learning. This aims to refine aesthetic appeal and text-image alignment without compromising the subject fidelity. We adapt the efficient MixGRPO framework [30], which confines stochastic optimization to a sliding window $S$. However, standard GRPO with its token-level policy ratios can exhibit instability, particularly when training MoE models due to expert routing fluctuations [71]. To mitigate this and better suit our MoE-LoRA architecture from Section˜4.2, we replace the GRPO objective with the more stable Group Sequence Policy Optimization (GSPO) [71]. Specifically, we replace the token-level policy ratio in Eq.˜3 with a sequence-level policy ratio $s_{i}(\theta)$ defined over the denoising steps within the sliding window $S$:
@@
This sequence-level ratio reflects the overall policy shift for the entire sequence within the optimization window, leading to more stable gradients. The advantage score $A_{i}$ is calculated as in Eq.˜3, but using our composite reward from Eq.˜11. The final optimization objective is thus:
$J(\theta)=\mathbb{E}_{x\sim\pi_{\theta_{\text{old}}}(\cdot|c,\mathcal{Z})}$ $\left[\frac{1}{N}\sum_{i=1}^{N}\min\Big(s_{i}(\theta)A_{i},\right$. (10)
$\left.clip(s_{i}(\theta),1-\beta,1+\beta)A_{i}\Big)\right].$
@@
where $R_{text}$ is the text alignment reward from a pre-trained CLIP model, $R_{aes}$ is an aesthetic reward from a predictor like HPSv2 [60], $R_{id}$ is used to evaluate subject fidelity, and $w_{id},w_{text},w_{aes}$ are their corresponding weights, . To accurately measure the subject fidelity of multi-subject generation results, we built the Multi-ID Alignment Reward using the Hungarian matching algorithm. For humans, we first employ a face detector [7] to extract facial embeddings from each reference image. We then apply the same detector to the generated image to identify all faces and extract their embeddings. Then we construct a pairwise similarity matrix $C$ where $C_{ij}$ is the cosine similarity between the embedding of the $i$-th reference face and the $j$-th detected face. The Hungarian algorithm is then used to solve the assignment problem by finding an assignment matrix $X\in\{0,1\}^{N_{ref}\times N_{gen}}$ that maximizes the total similarity:
@@
where $N_{ref}$ and $N_{gen}$ are the number of reference and generated faces, respectively. These ensure each face is matched at most once, preventing reward hacking, stopping the model from using attribute leakage to generate multiple “average faces" for an unearned high reward. For object subjects, each reference object is pre-annotated with a text prompt. We leverage Florence-2 [61] and SAM2 [47] to locate the corresponding object in the generated image. Then we compute the cosine similarity between the DINOv2 [44] embeddings of the segmented region and the reference object. We provide the complete pseudocode in the Appendix.
@@ 5.1 Experimental Setup
CLIP-T Face-Sim DINO-I CLIP-I AES CLIP-T DINO-I CLIP-I AES AVG
@@
Table 1: Quantitative comparison with SoTA methods on the multi-human and multi-object generation. The best results are in bold, and the second-best are underlined. Our method outperforms others, especially in subject fidelity, and achieves the highest overall score.
@@
Implementation Details. To achieve accurate multi-subject customized generation, our training process is decoupled into two stages. Following [59], we set the resolution of generated images to $512\times 512$ , the resolution of reference images to $320\times 320$, and each LoRA module’s rank to $r=512$. For the MoE-LoRA applied to the FFN layers, we configure it with 4 experts and activate 1 expert per forward pass. For reinforcement learning, we configure with a sampling step 16, a window size of $w=2$, a shift interval of $\tau=50$, and a window stride of $s=1$. More detailed settings can be found in the Appendix.
@@
Datasets and BenchMark. We constructed separate datasets for multi-human and multi-object generation. Due to the scarcity of public multi-human customization datasets with adequate annotations, we designed a data collection pipeline to curate a new dataset from OpenHuman-Vid datasets [li2025openhumanvid]. We supplemented the details in the Appendix. A total of 200k pairs of cross-pair data are obtained for multi-human generation training. For the reinforcement learning, we curated a face collection of 80 celebrities and 80 non-celebrities. We paired the faces in the collection and used Qwen2.5-VL [1] to generate a prompt for each pair, resulting in 12,720 data points. We randomly selected 1,000 of these as our benchmark, and the remaining data is used as reinforcement learning training data. For multi-object customization, we use the public MUSAR-Gen [17] dataset as our foundation. We use Florence-2 and SAM to obtain the detailed positions of reference objects. To ensure segmentation quality, we further employ Qwen2.5-VL to filter the results. We split 1,000 samples from this dataset that are not seen during training as a benchmark.
@@
Evaluation metrics. Following the  [28, 43], we evaluate generated image quality using standard metrics. We calculate the cosine similarity between the CLIP [46] embeddings of the prompt and the image (CLIP-T) to evaluate text fidelity. We also use HPSv2 for aesthetic scoring. For subject fidelity, we employ cosine similarity measures between generated images and reference images within CLIP and DINO [67] spaces, referred to as CLIP-I and DINO-I scores, respectively. Note that to accurately calculate CLIP-I and DINO-I, we first use a face detector [6] or Florence-2 to accurately locate the position of subjects in the generated image, and then calculate them. Additionally, for multi-human generation, we incorporate Face Similarity (Face-Sim) [6] and the Hungarian algorithm, enabling them to assess subject fidelity more accurately.
@@
Figure 4: Qualitative comparison with existing methods on the multi-subject generation (Zoom in for best visual comparison). Our method significantly improves subject fidelity while maintaining good text alignment and aesthetics.
@@ 5.2 Quantitative Comparison
We conducted comprehensive comparisons with existing methods in both multi-human and multi-object customized generation. These methods include MS-Diffusion [56], MIP-Adapter [72], OmniGen [62], UNO [59], OmniGen2 [58], DreamO [43], and XVerse [3]. The results, as shown in Tab.˜1, demonstrate that our method achieves significant improvements over existing approaches, achieving the highest overall score. In the multi-human generation task, our model shows a commanding lead in subject fidelity, achieving the top score. Notably, the significant 28.3% relative improvement in the Face-Sim over the next-best method highlights our model’s ability to effectively distinguish between different subjects and preserve their individual detailed features. This is crucial for accurately and reliably accomplishing the highly sensitive and challenging task of multi-human customization. Concurrently, the model remains highly competitive in text-image alignment. This outstanding performance extends to the multi-object generation benchmark, where our method once again achieved top ranks in text alignment and thematic fidelity, and second place in aesthetics, validating its robustness and versatility. This demonstrates the effectiveness of our decoupling framework, which achieves a strong balance between subject fidelity and multi-dimensional human preferences. It’s worth noting that the decline in aesthetic metrics for multi-person customized generation is due to the quality of the training data. We’ve provided further details in Appendix. Competitive results achieved on high-quality, open-source, multi-object custom data further support this.
@@ 5.3 Qualitative Comparison
To better demonstrate the effectiveness of our method, we focused on the more challenging task of multi-human customized generation and conducted a qualitative comparison with existing methods.. As shown in the first two rows of Fig.˜4, methods trained directly with In-Context Learning, such as UNO, OmniGen, and OmniGen2, struggle with attribute confusion when generating subjects of the same gender, which degrades subject fidelity. In contrast, our method accurately preserves the unique features of each individual while maintaining high text alignment and image quality. Our method maintains strong subject fidelity even in interactive scenarios, as shown in the third row of Fig.˜4. This result validates the effectiveness of our framework. For multi-objects generation, as shown in the last two lines of Fig.˜4, our method also demonstrates good subject fidelity and generated image quality. Appendix provides further qualitative comparisons of multi-subject generation (including human and objects) and includes a User Study to demonstrate that our method captures more human preferences.
@@ 5.4 Ablation Studies
$\mathcal{L}_{attn}$ MoE-LoRA IPPO CLIP-T Face-Sim DINO-I CLIP-I AES Overall
0.2645 0.1474 0.5972 0.6489 0.2954 0.3907
@@
Table 2: Ablation study of our methods. $\mathcal{L}_{attn}$ is Eq.˜5 in Identity-Disentangled Attention Regularization, and IPPO is our Identity-Preserving Preference Optimization. The results demonstrate the effectiveness of each module.
Figure 5: Ablation study on the effectiveness of MoE-LoRA. MoE-LoRA expands the model capacity, enabling the model to better grasp the diversity of spatial layout.
We conduct detailed ablation studies to validate the effectiveness of each component of our decoupling framework. Since the multi-human customization evaluation metrics are more accurate, our quantitative ablation experiment is mainly based on multi-human customization. We use UNO [59], a Flux-based model trained with highly coupled objectives, as our baseline for comparison, and the detailed results are presented in Tab.˜2 and Fig.˜5.
Effect of the Identity-Disentangled Attention Regularization. Identity-Disentangled Attention Regularization aims to achieve high-level subject fidelity in post-training. As shown in Tab.˜2, the base model performs poorly in terms of subject fidelity, achieving a Face-Sim score of only 0.1474. Upon introducing our $\mathcal{L}_{attn}$ using single LoRA, we observe a dramatic improvement across all subject fidelity metrics. As shown on the left of Fig.˜2, our method successfully separates the attention of different subjects. However, it can be seen that $\mathcal{L}_{attn}$ leads to a decrease in text alignment. As shown in Fig.˜5, this inability of a single LoRA to learn complex spatial layouts caused by different prompts and subjects results in some spatial layouts failing to be generated correctly, e.g., unable to generate the scene of a teddy bear riding on a motorcycle.
Building on $\mathcal{L}_{attn}$, we integrate the MoE-LoRA architecture to expand model capacity, which yields further gains across subject fidelity while also improving text alignment. Furthermore, as shown in Fig.˜5, the combinatorial samples that could not be generated correctly before have been significantly improved. This demonstrates that utilizing MoE-LoRA to expand model capacity can effectively alleviate the challenges of diverse spatial layouts.
Effect of Identity-Preserving Preference Optimization. We introduce IPPO to align human preferences and ensure subject fidelity during post-training. As shown in  Tab.˜2, the introduction of IPPO significantly improves text alignment and aesthetic metrics, while further enhancing subject fidelity, demonstrating the effectiveness of this method.
@@ 6 Conclusion
In this paper, we address the coupled training objective of existing multi-subject generation methods, where a single loss fails to optimize for both subject fidelity and human preference. We introduce a framework that decouples this problem into two stages. First, a pre-training stage uses explicit positional supervision to fix attention bleeding and achieve high fidelity. Second, a post-training stage employs an identity-preserving reinforcement learning framework to align with human preferences while maintaining this fidelity. Our decoupled approach achieves state-of-the-art results in both fidelity and preference alignment, validating its effectiveness.
Supplementary Material
@@ Appendix A More Details for Identity-Preserving Preference Optimization
We provide a detailed algorithm execution flow for Identity-Preserving Preference Optimization Section˜4.3, as shown in Algorithm˜1.
@@
In this section, we provide more details on the hyperparameter settings and specific training details. As mentioned in the main text, we decouple the training process into two parts, and the details of each phase are as follows:
Subject Fidelity Focused Pre-training. As described in Section˜4.1, in this stage, we train the model to primarily ensure it can generate images with high subject fidelity. Following [59, 3], we adopt a progressive training approach, first pre-training on single-subject data, and then using this as a foundation to train on multi-subject data. 1. Single-Subject Pre-training. We first pre-train the model for 40,000 steps on an internal single-subject dataset to equip it with foundational subject customization capabilities. In this stage, only $\mathcal{L}_{diff}$ is used for supervision. We use the AdamW optimizer with a $3\times 10^{-5}$ learning rate and a weight decay of $1\times 10^{-2}$. We use 8 cards for training and set the batch size of each card to 6. 2. Multi-Subject Customization Training. Then we train two models for customized human generation and object generation, respectively. For multi-human customized generation, we decrease the learning rate to $1\times 10^{-5}$, set the loss weight $\lambda=0.3$, and introduce the attention loss $\mathcal{L}_{attn}$. This stage runs for 25,000 steps. We use 8 cards at this stage and set the batch size to 4. For multi-object customized generation, since the size of the dataset is smaller than multi-human datasets, we only train for 15,000 steps.
Multi-Dimensional Preference Alignment Post-training. As described in Section˜4.1, after completing the Subject Fidelity Focused Pre-training, in the second stage, we primarily use post-training to align it with multi-dimensional human preferences. We fine-tune the model using our proposed Identity-Preserving Preference Optimization. Following [30], we configure with a sampling step of 16, a window size of $w=2$, a shift interval of $\tau=50$, and a window stride of $s=1$. This stage consists of 300 steps. For multi-human customization, the reward weights are set to $w_{id}=0.5$, $w_{text}=1.4$, and $w_{aes}=0.7$. For multi-object customization, we adjust the subject fidelity weight to $w_{id}=1.0$, while keeping $w_{text}=1.4$ and $w_{aes}=0.7$. In this stage, we used 16 cards for training and set the batch size of each card to 1.
@@
Due to the scarcity of public multi-human customization datasets with adequate annotations, we designed a comprehensive and automated data preparation process to extract training data from OpenHumanVid [li2025openhumanvid] video datasets, as shown in Fig.˜1. This pipeline processes raw video clips to generate structured training samples, each containing a target image with multiple subjects, corresponding identity reference images, segmentation masks, and a detailed textual description. The entire workflow ensures subject fidelity, high image quality, and rich annotation. Specifically, we sample video clips featuring two individuals from a large database. The process begins with frame selection and subject localization. For each input video, we sample an initial frame as the source for reference images and a middle frame as the target scene. We first employ a YOLO-Pose [40] to obtain initial bounding boxes and keypoint information for each person. Following localization, we leverage the Segment Anything Model (SAM) [47] to generate high-fidelity segmentation masks for each individual, effectively isolating them from the background. To refine this output, only the largest connected component of the mask is retained. A critical subsequent step is ensuring subject fidelity across frames. We apply a face detection model [6] to the segmented portraits to locate facial regions, and then use a face recognition model [6] to extract a normalized feature embedding for each face. By computing the cosine similarity between embeddings from the reference and target frames, we enforce a stringent threshold to discard pairs where identity cannot be confidently verified. Once a pair of frames passes this verification, the pipeline generates the final training sample. The segmented portraits and cropped faces from the initial frame are saved as the reference =images. The target frame is cropped into a square as target image, with its corresponding body and face masks preserved. Finally, a powerful vision-language model, Qwen2.5-VL [1], is prompted with the reference images and the target image to produce a rich text prompt of the entire scene, ensuring descriptive consistency for each subject.
@@
Figure 1: Data processing pipeline for customized multi-human image generation.
Table 1: Analysis of Data Quality and Preference Alignment. We compare the original UNO with a version retrained on our dataset (UNO). Training with UNO on our data also caused the aesthetic metrics to drop to levels similar to those before we used IPPO.
Methods CLIP-T $\uparrow$ Face-Sim $\uparrow$ DINO-I $\uparrow$ CLIP-I $\uparrow$ AES $\uparrow$
UNO (Official) 0.2645 0.1474 0.5972 0.6489 0.2954
[…]
Figure 2: Aesthetic Score Distribution of Training Data. The histogram illustrates the frequency of HPS v2 scores within our training dataset. The distribution is centered around a mean of 0.2552.
Appendix D Impact Analysis of Dataset Quality
To investigate whether the variance in aesthetic metrics between our method and certain baselines for multi-human generation task (e.g., UNO [59]) stems from algorithmic limitations or discrepancies in training data quality, we conducted a detailed statistical analysis and ablation study. It is important to note that many existing methods utilize private, high-quality datasets, whereas our model is trained on a subset curated from open-source video data, which inevitably contains frames with motion blur or lower aesthetic appeal.
Data Aesthetic Distribution. We first analyzed the aesthetic quality of our training dataset using the HPS v2 scoring model. As illustrated in Fig. 3, the aesthetic scores of our training samples follow a normal distribution with a mean of $0.2552$ and a median of $0.2556$. This relatively low baseline suggests that the model’s "upper bound" for aesthetics is naturally constrained by the training data when using standard supervised learning.
Baseline Retraining and Comparison. To verify this hypothesis, we conducted ablation study using UNO [59] as a baseline. We retrained the UNO model on our dataset (denoted as UNO) and compared it with the official UNO release and our method. As shown in Tab. 1, when UNO is retrained on our data, its aesthetic score (AES) drops significantly from $0.2954$ (Official) to $0.2656$. Crucially, this result ($0.2656$) is highly consistent with the performance of our method before the post-training stage (Ours w/o IPPO, $0.2661$). This observation supports two key conclusions:
  • The decline in aesthetic quality is primarily attributable to the training data rather than the model architecture.
  • Even with suboptimal data, our full method (Ours Full) successfully recovers the aesthetic quality to $0.2915$ via the proposed Identity-Preserving Preference Optimization (IPPO).
This demonstrates that our two-stage decoupled framework effectively aligns the model with human aesthetic preferences, overcoming the limitations of the underlying training data quality.
Appendix E More Details for Our Benchmark.